OpenStudy (anonymous):

Zach ran t times around the track at school, which is one-fourth mile. He ran a total of seven-eighths mile. Solve for t.

OpenStudy (johnweldon1993):

Well each time he runs around it = 1/4 of a mile....If he ran a total of 7/8 of a mile....he ran... 1/4t = 7/8 Multiply both sides by 4 to isolate t \[t = \frac{7}{8} \times 4\] t = ?

OpenStudy (anonymous):

two-sevenths seven thirty-seconds three and one-half four and four-sevenths

OpenStudy (johnweldon1993):

Right I see that the answer is there....just solve the equation above to find it...

OpenStudy (anonymous):

I GOT THE THIRD ANSWER

OpenStudy (johnweldon1993):

That is correct!